OpenStudy (anonymous):

Find each value if f(x)= 3x-5 and g(x)=x^2=x 1.) f(-3) 2.) f(2/3) 3.) g(5n)

hero (hero):

f(-3) = (3(-3)-5) = -14 f(2/3) = (2-5) = -3 g(5n) = ? You wrote g(x) = x^2 = x <----- I'm positive that isn't correct as written

hero (hero):

Do you see your mistake?

hero (hero):

repost g(x) correctly if possible

hero (hero):

Sometimes it's hard to notice your mistake I suppose

OpenStudy (anonymous):

yes but for some reason it has g(5n) and thank~you because I'm not good at math

hero (hero):

Yeah, I figured you won't notice your own mistake: I'll give you some choices... Is it: g(x) = x^2 = x g(x) = x^2 + x g(x) = x^2 - x Which one is it?

OpenStudy (anonymous):

neither I gt the answer thanks for helping me

hero (hero):

I was referring to the original function you posted. It has to be at least one of those because you posted it. Except the one you posted is definitely not the correct g(x) because x^2 does not equal x unless x equals zero or one.

OpenStudy (anonymous):

g(5n) = ? 25n^-5n this is the answer

hero (hero):

Okay, so I believe your original g(x) = x^2 - x which means.... g(5n) = (5n)^2 - 5n = 25n^2 - 5n

OpenStudy (anonymous):

yes your correct